OpenStudy (linstead_cpd_freak):

What is the solution to the equation? 6 + 2s – 8s = 18 A. s = –4 B. s = –2 C. s = 2 D. s = 4

OpenStudy (linstead_cpd_freak):

I think maybe A or B

OpenStudy (booklover40):

Hang on...

Mehek (mehek14):

well first combine 2s - 8s = ?

OpenStudy (booklover40):

can i tell you the answer?

OpenStudy (linstead_cpd_freak):

-6s

Mehek (mehek14):

@booklover40 direct answers aren't allowed :/ You have to guide the person

OpenStudy (booklover40):

I was going to give it and explain,but sorry.

Mehek (mehek14):

ok so 6 - 6s = 18 subtract 6 18 - 6 = ?

OpenStudy (linstead_cpd_freak):

12

OpenStudy (linstead_cpd_freak):

@Mehek14 What next?

Mehek (mehek14):

alright so we have -6s = 12 divide by -6 s = 12/-6 = ?

OpenStudy (linstead_cpd_freak):

-2

Mehek (mehek14):

so s = -2

OpenStudy (linstead_cpd_freak):

Thank you

Mehek (mehek14):

You're welcome :)